esha1963:

Andrea is designing the seating arrangement for a concert in her local park. To give everyone a good view, each row must have 4 more seats than the row before it, and the 1st row can only have 10 seats. Explain to Andrea how to create an equation to predict the number of seats in any row. Then, using your equation, show your work to determine the number of seats in row 20.

esha1963:

@Hero

esha1963:

@darkknight

esha1963:

@silvernight269

darkknight:

So we can set this up as a linear function. Where 4 is our slope (because it moves up 4, every row), and we also know that the first row has 10. So we can have a coordinate point as (1,10). So do you know point-slope form?

darkknight:

Point-slope form is in this form: Where x1 and y1 are the coordinate points of a point on the line. \[y-(Ycoor) = m(x-(Xcoor))\]

esha1963:

i think i figured it out

darkknight:

So we can plug in a few points. Such as 1 for the Xcoor, 10 for the Ycoor. And 4 for m. \[y-10=4(x-1)\] And then make it a slope-intercept form

esha1963:

can you confirm if my answer is correct?

darkknight:

Then you plug in 20 for x.

darkknight:

What did you get?

esha1963:

i used s+4(n-1) where s is the number of seats in the first row. n is the row #

esha1963:

so 10+4(20-1) i got 86 as the answer

darkknight:

A different way to do it. But correct nevertheless! Good Job

esha1963:

THANKS

esha1963:

i have another one like that

darkknight:

Sure, new post though. Close this one.

esha1963:

ok